OpenStudy (anonymous):

when you find factors for a number and you are trying to find a perfect sqaure for the radical and if they have more than pne do you use the biggest or smallest perfect sqaure????? please help mi reward is a medal.....

OpenStudy (amistre64):

any perfect square will help reduce the work; the largest perfect square will reduce the work the fastest

OpenStudy (amistre64):

if you have many perfect squares, you can take them all out ...

OpenStudy (amistre64):

if you have an example youd like to work on, that might help make the point

OpenStudy (anonymous):

\[\sqrt{72}\]

OpenStudy (amistre64):

and 72 factors into .... any perfect squares?

OpenStudy (amistre64):

9*8 9*4*2, or simply 36*2 we can either take out (3*2) for the perfect squares of 9 and 4 or we can just take out a 6 for the perfect square of 36, either way, we get 6 sqrt2

OpenStudy (amistre64):

\[\sqrt{72}\] \[\sqrt{9*4*2}\] \[\sqrt{9}\sqrt{4}\sqrt{2}\] \[3\sqrt{4}\sqrt{2}\] \[3*2\sqrt{2}\] \[6\sqrt{2}\]

OpenStudy (anonymous):

thats what i was thinkingbut if i use 36 and 2 it would comm out faster...

OpenStudy (amistre64):

faster yes, but if you cant see it as 36*2 to start with, then the smaller factors still produce the same resutls :)